OpenStudy (anonymous):

Please Help if sin tetha=12/13 and cos thetha <0, what is tan tetha??

OpenStudy (kamille):

well, do you know how to find cos thetha?

OpenStudy (anonymous):

sin is positive and cos is negative. That is Q II. sin t = opp / hyp = 12 / 13 therefore adj side is 5. Tan t = opp/adj = -12/5 is that correct??

OpenStudy (kamille):

Yes:)